OpenStudy (anonymous):

Solve the equation, if possible. 18 cos 3x-9=0

OpenStudy (michele_laino):

please, add 9 to both sides of your equation, subsequently divide the resultant equation by 18, please write your answer

OpenStudy (anonymous):

cos3x=1/2

OpenStudy (michele_laino):

that's right!, now you know that 3x must equal to \[\pi/3 +2*k*\pi \] where k is an integer with sign, please try to calculate x and write your answer

OpenStudy (anonymous):

How did you get the π/3+2∗k∗π out of my answer??

OpenStudy (michele_laino):

because when \[\cos \alpha =1/2\] then \[\alpha =60°\] nevertheless 60° is not the unique solution, there are infinite solutions, which you can get adding to 60° the quantity k*360°, where \[k=\pm 1, \pm2, \pm3, ....\]

OpenStudy (michele_laino):